OpenStudy (anonymous):

solve ln (a +10) =2 ln (a-2)

OpenStudy (anonymous):

Do you know the log rules?

OpenStudy (anonymous):

The easiest way to go about it is using the rule. ln (a+10) = ln (a-2)^2

OpenStudy (anonymous):

You are allowed to bring the 2 from in front of the ln as an exponent. Now you can drop the ln an solve it as (a+10) = (a-2)^2. Can you handle it from here?

OpenStudy (anonymous):

ok i will try

OpenStudy (anonymous):

Just tag me if you need further help

OpenStudy (anonymous):

ok..thank you so much :D